OpenStudy (anonymous):

find the value of c if the limit of x^2+cx+4devided by x-5 exists as x approaches 5

OpenStudy (anonymous):

Take \[c= -\frac {29}4 \] and notice that \[x^2-\frac{29 x}{5}+4 =\frac{1}{5} (x-5) (5 x-4) \] \[ \frac {\frac{1}{5} (x-5) (5 x-4)} {x-5}= \frac 1 5 (5x-4) \] which tends to \[ \frac {21}5\] when x tend to 5

OpenStudy (anonymous):

Try to figure out, how one can find \[ c= - \frac {29}4\]